One angle of a triangle is three times another. The third angle is 45 degrees greater than the smallest. Find the measure of the smallest angle.
Thanks

Answer by stanbon(75887)

One angle of a triangle is three times another. The third angle is 45 degrees greater than the smallest. Find the measure of the smallest angle.
-----------------------
A + b + c = 180
a = 3b
c = b + 45
----------------------
Rearrange:
a + b + c = 180
a -3b + 0 = 0
0 + b - c = -45
-----------------------
Use any method you know to get:
a = 81
b = 27
c = 72
=============
Cheers,
Stan H.